A random sample of 16 female college-aged dancers showed a sample mean height 64.8 of inches and a sample standard deviation of 1.7 inches.
1.Find a 95% confidence interval for the population mean height of dancers and interpret the interval.
2. Find a 99% confidence interval for the population mean height and interpret the interval.
3. Which interval is wider and why?
Thank you!!

Answer by ewatrrr(24785)

Sample: n = 16, mean= 64.8 , sd = 1.7
ME = 1.96(1.7/sqrt(16) = .8 (64, 65.6)
ME = 2.576(1.7/sqrt(16) = 1 (63. 65.8) ***wider
Including 99% rather than just 95%
.........
for Your reference:
CI
90% z =1.645
92% z = 1.751
95% z = 1.96
98% z = 2.326
99% z = 2.576